Complete the following probability distribution table and then calculate the stated probabilities. HINT [See Quick Example 5.]

Outcome a b c d e
Probability 0.1 0.05 0.4 0.05

(a)    P({a, c, e})
P({a, c, e}) =  

(b)    P(EF), where E = {a, c, e} and F = {b, c, e}
P(EF) =  

(c)    P(E'), where E is as in part (b)
P(E') =  

(d)    P(EF ), where E and F are as in part (b)
P(EF) =  
Answer #1

P(e)=1-(P(a)+P(b)+P(c)+P(d))

=1-(0.1+0.05+0.4+0.05)=0.4

a)

P({a,c,e}) = P(a)+P(c)+P(e)

=0.1+0.4+0.4=0.9

b)

P(EF) = P(a,b,c,e)

P(a)+P(b)+P(c)+P(e) = 0.6

c)

P(E')=P(b,d)

=0.05+0.05=0.1

d)

P(EF) = P({c,e})

=0.4+0.4 = 0.8